# AnyAnomaly: Zero-Shot Customizable Video Anomaly Detection with LVLM

### v1

@@ Abstract
Video anomaly detection (VAD) is crucial for video analysis and surveillance in computer vision. However, existing VAD models rely on learned normal patterns, which makes them difficult to apply to diverse environments. Consequently, users should retrain models or develop separate AI models for new environments, which requires expertise in machine learning, high-performance hardware, and extensive data collection, limiting the practical usability of VAD. To address these challenges, this study proposes customizable video anomaly detection (C-VAD) technique and the AnyAnomaly model. C-VAD considers user-defined text as an abnormal event and detects frames containing a specified event in a video. We effectively implemented AnyAnomaly using a context-aware visual question answering without fine-tuning the large vision language model. To validate the effectiveness of the proposed model, we constructed C-VAD datasets and demonstrated the superiority of AnyAnomaly. Furthermore, our approach showed competitive performance on VAD benchmark datasets, achieving state-of-the-art results on the UBnormal dataset and outperforming other methods in generalization across all datasets. Our code is available online at github.com/SkiddieAhn/Paper-AnyAnomaly.
@@ 1 Introduction
Video anomaly detection (VAD) aims to detect abnormal events in video streams. Abnormal events include the actions of objects that are inappropriate for the environment (e.g., climbing over a fence) or objects with unusual appearances (e.g., a bicycle on a walkway). However, abnormal events are rare and diverse, making it difficult to construct large-scale datasets for VAD. Therefore, the VAD is recognized as a highly challenging problem.
To overcome these limitations, previous studies primarily used one-class classification (OCC) methods that learn only from normal data. In the OCC approach, the model learns normal patterns and classifies the cases that deviate from them as abnormal. Representative OCC-based VAD methods include classification-[27, 35, 14], distance-[26, 2, 24], and prediction-based methods [15, 33, 10], all of which have demonstrated excellent performance in VAD tasks.
@@
An effective method to implement zero-shot C-VAD is to leverage large vision language models (LVLMs). Recently, LVLMs have demonstrated outstanding generalization performance in visual text analysis. By leveraging this capability, C-VAD can be performed effectively across various environments. The most intuitive method involves performing visual question answering (VQA) [4] on each frame to estimate the anomaly score. For instance, one could provide the model with the prompt: “Return a value between 0 (no) and 1 (yes) indicating how well the input image represents the text provided by the user”. This was used as the baseline model. However, through experiments, we observed the following limitations of the baseline model: 1 Due to the large computational cost of LVLMs, the latency is high. 2 Difficulty in analyzing specific objects due to the characteristics of surveillance videos, such as foreground-background imbalance and object congestion. 3 Difficulty in detecting action-related anomalies because of the inability to utilize temporal information.
To overcome these limitations, we designed an AnyAnomaly model with the structure shown in Figure 2. First, to reduce the latency, we adopted a segment-level approach that groups consecutive frames into a single segment for processing. For this purpose, we introduced a key frames selection module (KSM) that selects key frames representing the segment and performed VQA per segment. Second, instead of performing simple image-text matching, we introduced a context-aware VQA approach to enable a deeper understanding of the scene. To this end, we additionally utilized two types of information: position context, $PC$ and temporal context, $TC$. $PC$ is a context that emphasizes important locations within a frame, enhancing the object analysis capability of the LVLM. $TC$ is a context that structures scene changes over time into a grid format, improving the action analysis capability of the LVLM. Notably, the proposed KSM and context generation modules operate in a training-free manner, allowing for easy application of C-VAD without additional data training.
To evaluate the performance of C-VAD, we classified existing VAD benchmark datasets based on anomaly types to create the C-VAD datasets. Through this process, we demonstrated the superiority of AnyAnomaly. Furthermore, despite being a zero-shot approach, AnyAnomaly achieved competitive performance on VAD datasets compared to traditional OCC-based VAD models. It achieved state-of-the-art (SOTA) results on the UBnormal dataset [1] and showed superior generalization across all datasets. The proposed approach is expected to be an effective solution for deploying VAD technology in real-world applications. The contributions of this study are as follows:
@@
  • AnyAnomaly achieves SOTA performance on UBnormal dataset and outperforms other methods in generalization across all datasets.
@@ 2 Related Work
Video Anomaly Detection. Most VAD models adopt the OCC approach to detect anomalies by learning normal patterns. Among them, prediction-based methods train models to predict future or past frames based on normal frames, assuming that abnormal frames exhibit larger prediction errors. Liu [15] proposed a method that utilizes FlowNet [11] and GANs [20] to predict the $t+1$-th frame given $t$ input frames. Yang [33] introduced an approach that selects key frames from $t$ input frames to predict the entire sequence. However, because the definitions of normal and abnormal patterns may vary depending on the environment, the OCC approach, which relies on learned normal patterns, has a limited generalization performance. To mitigate this limitation, recent studies have explored cross-domain VAD (xVAD). Notably, zxVAD [3] enhances the adaptability to new environments by synthesizing abnormal patterns using the cut mix technique on images from auxiliary datasets. However, these approaches depend on fixed data transformations, making it difficult to fully capture the diverse abnormal patterns that may occur in real-world scenarios. Therefore, we propose a novel VAD method that uses textual information to dynamically detect abnormal patterns that vary depending on the environment.
Large Vision Language Models. Large language models have primarily been used in natural language processing; however, they have recently been applied to multimodal tasks such as image captioning and VQA. For example, MiniGPT-4 [38] processes multimodal inputs by connecting a pre-trained vision encoder to the Vicuna [7] model through a linear layer. Recent LVLMs have employed novel visual encoding techniques to better understand images. Chat-UniVi [13] generates dynamic tokens for images, thereby reducing unnecessary information and effectively extracting key visual features. This model enables flexible analysis by applying dynamic tokens across various resolutions. MiniCPM-V [34] applies the best partition technique according to the image resolution and generates tokens optimized for each segment, thereby improving the memory efficiency. However, despite the advancements in LVLMs, they are trained for general purposes, making their direct application in VAD challenging. To handle VAD effectively, it is essential to consider the characteristics of surveillance videos and leverage temporal information. Therefore, we propose a training-free approach to minimize the domain gap between the LVLMs and VAD tasks.
@@ 3.2 Key frames Selection Module
Figure 3(a) shows KSM, a key component of the segment-level approach. For this purpose, we selected four frames representing the segment as $K$ and utilized the CLIP [22] model, which was trained to match the images and text.
@@
The index of the representative frame $\hat{k}$, denoted as $\hat{i}$, is used to select the other key frames. We divide the segment into four groups of equal size and select the $\hat{i}\bmod\frac{N}{4}$-th frame from each group. For example, when $N=8$ and $\hat{i}=4$, the 00-th frame from each group is selected and the final set is $K=\{s_{0},s_{2},s_{4},s_{6}\}$. This process is defined as follows:
@@ 3.3 Context Generation
$PC$ and $TC$ are key elements of context-aware VQA, serving as additional information that complements the input image. $PC$ enhances the object analysis capability of the LVLM and is generated through WinCLIP-based attention (WA). $TC$ strengthens the action analysis ability of the LVLM and is created through grid image generation (GIG).
WinCLIP-based Attention. Figure 3(b) illustrates the WA method. We emphasize the regions related to $X$ at $\hat{k}$ based on WinCLIP, as proposed by Jeong [12]. First, $\hat{k}$ is divided into multiple windows, and individual embeddings are generated from each window using $E_{I}$. For example, when the image size is $240\times 240$, it is divided into 25 windows of size $48\times 48$, and the embeddings of each window are collected to create a small-scale window embedding map $W^{s}\in\mathbb{R}^{25\times D}$. By adjusting the window size, a middle-scale window embedding map $W^{m}$ and large-scale window embedding map $W^{l}$ are also generated, and the similarity between these embedding maps and the text embedding $z\in\mathbb{R}^{D}$ is calculated. The final similarity map $M$ is generated by averaging the similarities calculated on three scales:
@@
We combined the template proposed by Jeong [12] with $X$ and passed it through $E_{T}$ to generate $z$. Finally, we multiplied $M$ and $\hat{k}$ to create $PC$:
@@
Grid Image Generation. Figure 3(c) illustrates the GIG method, which comprises two stages. In the multiscale grid generation stage, $K$ is used to create grid images at different scales. Similar to the process described in WA, each frame of $K$ is divided into multiple windows, and the windows at the same position are connected in a 2$\times$2 grid format to create a single grid image. This process is defined as follows:
@@ 3.4 Anomaly Detection
Prompt Design. Figure 4 illustrates the proposed prompt $P$. The prompt comprises three main components: ‘task’, ‘consideration’, and ‘output’. First, ‘task’ defines the operation the LVLM should perform, specifically evaluating whether $X$ is present in the image. Next, ‘consideration’ specifies factors to be taken into account during evaluation, while ‘output’ defines the format for presenting the evaluation results. To leverage the chain-of-thought [30] effect, we instructed the model to provide brief reasoning along with the anomaly score, rounded to one decimal place. When conducting VQA using $TC$, an additional element, ‘context’, is inserted between ‘task’ and ‘consideration’ in the prompt. This context element conveys the meaning of the rows and columns of $TC$ to the LVLM. We define the modified prompt as $P^{*}$. A comparative analysis of prompts is presented in supplementary materials.
Anomaly Scoring. The context serves as supplementary information to the image. However, because the LVLM accepts only a single image as input, it is challenging to utilize both the original and additional information simultaneously. To address this issue, we adopt a late fusion approach. Specifically, $\hat{k}$, $PC$, and $TC$ were used as the image inputs for the LVLM. The LVLM returns an anomaly score for each input, and these three scores are combined to compute the final $ascore$:
@@
Here, $\gamma$ is a hyperparameter that adjusts the proportion of context reflected in $ascore$. A performance comparison experiment based on the hyperparameter tuning is presented in supplementary materials.
Consequently, even if the abnormal frame $\hat{k}$ receives a low score, the final anomaly score will be high if the additional information, $PC$ or $TC$, is assigned a high score. This enabled accurate anomaly detection.
@@ 4.1 Datasets
Figure 5 illustrates the composition of the VAD and proposed C-VAD datasets. In conventional VAD datasets, videos are not categorized by an abnormal class type. In contrast, the proposed C-VAD datasets are organized by abnormal event type, with videos classified as positive or negative based on the presence of each abnormality. This categorization enables a precise evaluation of detection performance for specific types of abnormalities (e.g., bicycle). In this study, we validated the effectiveness of the proposed method on three VAD datasets: CUHK Avenue (Ave) [16], ShanghaiTech Campus (ShT) [18], and UBnormal (UB) [1] as well as two C-VAD datasets: Customizable-ShT (C-ShT) and Customizable-Ave (C-Ave). Further details of the datasets are provided in supplementary materials.
@@ 4.2 Evaluation Criteria
Table 1: Performance comparison on C-ShT dataset
@@ 4.3 Results
When only KSM was applied to the baseline, the execution time decreased in proportion to the segment length, whereas the average performance remained similar to that of the baseline. This is because the CLIP effectively selects representative frames for each segment, thereby compensating for the loss of temporal information. However, because it does not fully capture fine-grained spatio-temporal details, its performance significantly decreases for certain classes.
Therefore, we address these issues using the proposed contextual information. First, using $PC$ resulted in performance improvements of 5.64% and 3.62% compared to the KSM, as the LVLM focused on analyzing objects related to $X$. Additionally, applying $TC$ led to performance improvements of 8.38% and 14.43% over the KSM, respectively, with particularly notable enhancements observed in the action class. This indicates that utilizing the temporal information provided in the grid image is essential for action analysis.
@@ 4.4 Qualitative Analysis
Figure 6: Anomaly score comparison and context visualization
@@ 4.5 Ablation Study
Segment length and FPS. Table 3 presents the performance comparison and FPS based on different segment lengths. The baseline segment length was set to 1. It was observed that deriving anomaly scores at the segment level yields superior performance compared to the baseline, which relies on a single frame. The highest AUC performance was achieved when the segment length is set to 24, reaching 85.72% and 90.27% for C-ShT and C-Ave, respectively. However, excessively long segment length introduces irrelevant information into the temporal context, leading to a decrease in accuracy. Furthermore, performing VAD at the segment level resulted in a 594% improvement in the FPS compared with the baseline.
Table 3: Comparison on segment length
Segment length C-ShT C-Ave FPS
Baseline 78.01 79.43 0.96
8 83.83 83.96 2.67
16 83.45 87.45 4.49
24 85.72 90.27 6.67
32 82.50 85.94 8.45
@@
Window Size. Table 5 presents the experimental results based on the window sizes used in $PC$ and $TC$. For the action classes, the best performance was achieved with the large window size in C-ShT and the middle window size in C-Ave. This indicates that middle or large window sizes are more effective in capturing temporal movements and interactions between multiple objects. For appearance classes, the optimal performance was observed with the small window size in C-ShT and the middle window size in C-Ave, suggesting that the appropriate window size varies depending on the dataset owing to differences in object sizes. To enhance the generalization performance of the model, we adopted an approach that utilized all three window sizes and found that incorporating them yielded the best overall performance.
@@
Method Venue Training Ave ShT UB
AMMC-Net[6] AAAI 21 86.6 73.7 -
STEAL-Net[5] ICCV 21 87.1 73.7 -
MPN[19] CVPR 21 89.5 73.8 -
Zhong et al.[36] PR 22 88.9 70.7 -
DLAN-AC[32] ECCV 22 89.9 74.7 -
UBnormal[1] CVPR 22 - - 68.5
FPDM[31] ICCV 23 90.1 78.6 62.7
SLM[28] ICCV 23 90.9 78.8 -
USTN-DSC[33] CVPR 23 89.9 73.8 -
MULDE[21] CVPR 24 - 81.3 72.8
AED-MAE[25] CVPR 24 91.3 79.1 58.5
MA-PDM[37] AAAI 25 91.3 79.2 63.4
AccI-VAD[24] TMLR 25 - - 66.8
AnyAnomaly* - 81.4 77.2 73.1
AnyAnomaly - 87.3 79.7 74.5
@@
STEAL-Net[5] 54.3 51.7
Jigsaw[29] 62.9 59.3
rGAN[17] 76.6 77.9*
MPN[19] 78.9 73.8*
zxVAD[3] 82.2 71.6*
Shibao et al.[8] 86.2 78.7
ZS CLIP[22] 62.3 60.9
ZS ImageBind[9] 64.5 61.3

### v4

@@ Abstract
Video anomaly detection (VAD) is crucial for video analysis and surveillance in computer vision. However, existing VAD models rely on learned normal patterns, which makes them difficult to apply to diverse environments. Consequently, users should retrain models or develop separate AI models for new environments, which requires expertise in machine learning, high-performance hardware, and extensive data collection, limiting the practical usability of VAD. To address these challenges, this study proposes customizable video anomaly detection (C-VAD) technique and the AnyAnomaly model. C-VAD considers user-defined text as an abnormal event and detects frames containing a specified event in a video. We effectively implemented AnyAnomaly using a context-aware visual question answering without fine-tuning the large vision language model. To validate the effectiveness of the proposed model, we constructed C-VAD datasets and demonstrated the superiority of AnyAnomaly. Furthermore, our approach showed competitive results on VAD benchmarks, achieving state-of-the-art performance on UBnormal and UCF-Crime and surpassing other methods in generalization across all datasets. Our code is available online at github.com/SkiddieAhn/Paper-AnyAnomaly.
@@ 1 Introduction
Video anomaly detection (VAD) aims to detect abnormal events in video streams. Abnormal events include the actions of objects that are inappropriate for the environment (e.g., climbing over a fence) or objects with unusual appearances (e.g., a bicycle on a walkway). However, abnormal events are rare and diverse, making it difficult to construct large-scale datasets for VAD. Therefore, VAD is recognized as a highly challenging problem.
To overcome these limitations, previous studies primarily used one-class classification (OCC) methods that learn only from normal data. In the OCC approach, the model learns normal patterns and classifies the cases that deviate from them as abnormal. Representative OCC methods include classification-[ruff2018deep, yi2020patch, lee2020multi], distance-[roth2022towards, ahn2024videopatchcore, reissattribute], and prediction-based methods [liu2018future, yang2023video, hong2024making, lee2025mdvad], all of which have demonstrated excellent performance in VAD tasks.
@@
An effective method to implement zero-shot C-VAD is to leverage large vision language models (LVLMs). Recently, LVLMs have demonstrated outstanding generalization performance in visual text analysis. By leveraging this capability, C-VAD can be performed effectively across various environments. The most intuitive method involves performing visual question answering (VQA) [antol2015vqa] on each frame to estimate the anomaly score. For instance, one could provide the model with the prompt: “Return a value between 0 (no) and 1 (yes) indicating how well the input image represents the text provided by the user”. This was used as the baseline model. However, through experiments, we observed the following limitations of the baseline model: 1 Due to the large computational cost of LVLMs, the latency is high. 2 Difficulty in analyzing specific objects due to the characteristics of surveillance videos, such as foreground-background imbalance and object congestion. 3 Difficulty in detecting action-related anomalies because of the inability to utilize temporal information.
To overcome these limitations, we designed an AnyAnomaly model with the structure shown in Figure 2. First, to reduce the latency, we adopted a segment-level approach that groups consecutive frames into a single segment for processing. For this purpose, we introduced a key frames selection module (KSM) that selects key frames representing the segment and performed VQA per segment. Second, instead of performing simple image-text matching, we introduced a context-aware VQA approach to enable a deeper understanding of the scene. To this end, we additionally utilized two types of information: position context, $PC$ and temporal context, $TC$. PCPC is a context that emphasizes important locations within a frame, enhancing the object analysis capability of the LVLM. $TC$ is a context that structures scene changes over time into a grid format, improving the action analysis capability of the LVLM. Notably, the proposed KSM and context generation modules operate in a training-free manner, allowing for easy application of C-VAD without additional training.
To evaluate the performance of C-VAD, we classified existing VAD benchmark datasets based on anomaly types to create the C-VAD datasets. Through this process, we demonstrated the superiority of AnyAnomaly. Furthermore, despite being a zero-shot approach, AnyAnomaly achieved competitive performance on VAD datasets compared to traditional OCC-based VAD models. Remarkably, It achieved state-of-the-art (SOTA) results on UBnormal [acsintoae2022ubnormal] and UCF-Crime [sultani2018real], and showed superior generalization across all datasets. The proposed approach is expected to be an effective solution for deploying VAD technology in real-world applications. The contributions of this study are as follows:
@@
  • AnyAnomaly achieves SOTA performance on UBnormal and UCF-Crime and outperforms other methods in generalization across all datasets.
@@ 2 Related Work
Video Anomaly Detection. Most VAD models adopt the OCC approach to detect anomalies by learning normal patterns. Among them, prediction-based methods train models to predict future or past frames based on normal frames, assuming that abnormal frames exhibit larger prediction errors. Liu [liu2018future] proposed a method that utilizes FlowNet [ilg2017flownet] and GANs [mao2017least] to predict the $t+1$-th frame given $t$ input frames. Yang [yang2023video] introduced an approach that selects key frames from $t$ input frames to predict the entire sequence. However, because the definitions of normal and abnormal patterns may vary depending on the environment, the OCC approach, which relies on learned normal patterns, has a limited generalization performance. To mitigate this limitation, recent studies have explored cross-domain VAD (xVAD). Both rGAN [lu2020few] and MPN [lv2021learning] adopted meta-learning approaches, enabling adaptation to a target domain from only a few training samples. zxVAD [aich2023cross] further improved adaptability to new environments by synthesizing abnormal patterns through the cut-mix technique applied to images from auxiliary datasets without explicit adaptation to the target domain. However, this approach depends on fixed data transformations, making it difficult to fully capture the diverse abnormal patterns that may occur in real-world scenarios. Therefore, we propose a novel VAD method that uses textual information to dynamically detect abnormal patterns that vary depending on the environment.
Large Vision Language Models. Large language models have primarily been used in natural language processing; however, they have recently been applied to multimodal tasks such as image captioning and VQA. For example, MiniGPT-4 [zhu2023minigpt] processes multimodal inputs by connecting a pre-trained vision encoder to the Vicuna [chiang2023vicuna] model through a linear layer. Recent LVLMs have employed novel visual encoding techniques to better understand images. Chat-UniVi [jin2024chat] generates dynamic tokens for images, thereby reducing unnecessary information and effectively extracting key visual features. This model enables flexible analysis by applying dynamic tokens across various resolutions. MiniCPM-V [yao2024minicpm] applies the best partition technique according to the image resolution and generates tokens optimized for each segment, thereby improving the memory efficiency. However, despite the advancements in LVLMs, they are trained for general purposes, making their direct application in VAD challenging. Therefore, we propose a training-free approach to minimize the domain gap between the LVLMs and VAD tasks.
@@ 3.1 Overview
The user-defined text $X$ refers to a natural language description of the anomaly the user wishes to detect. It can be a single word (e.g., “bicycle”), diverse events (e.g., “jumping-falling”), or a complex behavior (e.g., “driving outside the lane”). In the case of diverse events, each event keyword is processed individually as a single word.
@@ 3.2 Key frames Selection Module
Figure 3(a) shows KSM, a key component of the segment-level approach. For this purpose, we selected four frames representing the segment as $K$ and utilized the CLIP [radford2021learning] model, which was trained to match the images and text.
@@
The index of the representative frame $\hat{k}$, denoted as $\hat{i}$, is used to select the other key frames. We divide the segment into four groups of equal size and select the $\hat{i}\bmod\frac{N}{4}$-th frame from each group. For example, when $N=8$ and $\hat{i}=4$, the 0-th frame from each group is selected and the final set is $K=\{s_{0},s_{2},s_{4},s_{6}\}$. This process is defined as follows:
@@ 3.3 Context Generation
PCPC and $TC$ are key elements of context-aware VQA, serving as additional information that complements the input image. $PC$ enhances the object analysis capability of the LVLM and is generated through WinCLIP-based attention (WA). $TC$ strengthens the action analysis ability of the LVLM and is created through grid image generation (GIG).
WinCLIP-based Attention. Figure 3(b) illustrates the WA method. We emphasize the regions related to $X$ at $\hat{k}$ based on WinCLIP, as proposed by Jeong [jeong2023winclip]. First, $\hat{k}$ is divided into multiple windows, and individual embeddings are generated from each window using $E_{I}$. For example, when the image size is $240\times 240$, it is divided into 25 windows of size $48\times 48$, and the embeddings of each window are collected to create a small-scale window embedding map $W^{s}\in\mathbb{R}^{25\times D}$. By adjusting the window size, a middle-scale window embedding map $W^{m}$ and large-scale window embedding map $W^{l}$ are also generated, and the similarity between these embedding maps and the text embedding $z\in\mathbb{R}^{D}$ is calculated. The final similarity map $M$ is generated by averaging the similarities calculated on three scales:
@@
We combined the template proposed by Jeong [jeong2023winclip] with $X$ and passed it through $E_{T}$ to generate $z$. Finally, we multiplied $M$ and $\hat{k}$ to create $PC$:
@@
Grid Image Generation. Figure 3(c) illustrates the GIG method, which comprises two stages. In the multiscale grid generation stage, $K$ is used to create grid images at different scales. Similar to the process described in WA, each frame of $K$ is divided into multiple windows, and the windows at the same position are connected in a $\times2$ grid format to create a single grid image. This process is defined as follows:
@@ 3.4 Anomaly Detection
Prompt Design. Figure 4 illustrates the proposed prompt $P$. The prompt comprises three main components: ‘task’, ‘consideration’, and ‘output’. First, ‘task’ defines the operation the LVLM should perform, specifically evaluating whether $X$ is present in the image. Next, ‘consideration’ specifies factors to be taken into account during evaluation, while ‘output’ defines the format for presenting the evaluation results. To leverage the chain-of-thought [wei2022chain] effect, we instructed the model to provide brief reasoning along with the anomaly score, rounded to one decimal place. When conducting VQA using $TC$, an additional element, ‘context’, is inserted between ‘task’ and ‘consideration’ in the prompt. This context element conveys the meaning of the rows and columns of $TC$ to the LVLM. We define the modified prompt as $P^{*}$.
Anomaly Scoring. The context serves as supplementary information to the image. However, because the LVLM accepts only a single image as input, it is challenging to utilize both the original and additional information simultaneously. To address this issue, we adopt a late fusion approach. Specifically, $\hat{k}$, $PC$, and $TC$ were used as the image inputs for the LVLM. The LVLM returns an anomaly score for each input, and these three scores are combined to compute the final ascoreascore:
@@
Here, $\gamma$ is a hyperparameter that adjusts the proportion of context reflected in ascoreascore. A performance comparison experiment based on the hyperparameter tuning is presented in supplementary materials.
Consequently, even if the abnormal frame $\hat{k}$ receives a low score, the final anomaly score will be high if the additional information, $PC$ or $TC$, is assigned a high score. This enabled accurate anomaly detection. Finally, to create frame-level anomaly scores, we duplicated the ascoreascore for the length of each segment and then applied a temporal 1-D Gaussian filter for smoothing, following prior works [reissattribute, wang2022video].
@@ 4.1 Datasets
Figure 5 illustrates the composition of the VAD and proposed C-VAD datasets. In conventional VAD datasets, videos are not categorized by an abnormal class type. In contrast, the proposed C-VAD datasets are organized by abnormal event type, with videos classified as positive or negative based on the presence of each abnormality. This categorization enables a precise evaluation of detection performance for specific types of abnormalities (e.g., bicycle). In this study, we validated the effectiveness of the proposed method on four VAD datasets: CUHK Avenue (Ave) [lu2013abnormal], ShanghaiTech Campus (ShT) [luo2017revisit], UBnormal (UB) [acsintoae2022ubnormal], and UCF-Crime (UCF) [sultani2018real] as well as two C-VAD datasets: Customizable-ShT (C-ShT) and Customizable-Ave (C-Ave). Further details of the datasets are provided in supplementary materials.
@@ 4.2 Evaluation Criteria
Table 1: Performance comparison on C-ShT dataset. The best results are bolded. The second-best results are underlined.
@@ 4.3 Results
When only KSM was applied to the baseline, the execution time decreased in proportion to the segment length, whereas the average performance remained similar to that of the baseline. This is because the CLIP effectively selects representative frames for each segment, thereby compensating for the loss of temporal information. However, because it does not fully capture fine-grained spatio-temporal details, its performance significantly decreases for certain classes. Therefore, we address these issues using the proposed contextual information. First, using $PC$ resulted in performance improvements of 5.64% and 3.62% compared to the KSM, as the LVLM focused on analyzing objects related to $X$. Additionally, applying $TC$ led to performance improvements of 8.38% and 14.43% over the KSM, respectively, with particularly notable enhancements observed in the action class. This indicates that utilizing the temporal information provided in the grid image is essential for action analysis. Additional validations, such as FPS comparisons by segment length and performance evaluations across various LVLMs, are presented in supplementary materials.
@@ 4.4 Qualitative Analysis
Figure 6: Anomaly score comparison with context visualization
@@ 4.5 Ablation Study
Window Size. Table 4 presents the experimental results based on the window sizes used in $PC$ and $TC$. For the action classes, the best performance was achieved with the large window size in C-ShT and the middle window size in C-Ave. This indicates that middle or large window sizes are more effective in capturing temporal movements and interactions between multiple objects. For appearance classes, the optimal performance was observed with the small window size in C-ShT and the middle window size in C-Ave, suggesting that the appropriate window size varies depending on the dataset owing to differences in object sizes. To enhance the generalization performance of the model, we adopted an approach that utilized all three window sizes and found that incorporating them yielded the best overall performance. Further details on the window size are provided in the implementation details section of the supplementary material.
@@
Method Zero-shot Ave ShT UB UCF
AMMC-Net[cai2021appearance] 86.6 73.7 - -
STEAL-Net[astrid2021synthetic] 87.1 73.7 - -
MPN[lv2021learning] 89.5 73.8 - -
DLAN-AC[yang2022dynamic] 89.9 74.7 - -
UBnormal[acsintoae2022ubnormal] - - 68.5 -
FPDM[yan2023feature] 90.1 78.6 62.7 74.7
SLM[shi2023video] 90.9 78.8 - -
USTN-DSC[yang2023video] 89.9 73.8 - -
AnomalyRuler[Yang2024FollowTR] 89.7 85.2 71.9 -
MULDE[micorek2024mulde] - 81.3 72.8 78.5
AED-MAE[ristea2024self] 91.3 79.1 58.5 -
MA-PDM[zhou2024video] 91.3 79.2 63.4 -
AccI-VAD[reissattribute] - 76.2 66.8 60.3
AnyAnomaly* 81.4 77.2 73.1 77.8
AnyAnomaly 87.3 79.7 74.5 80.7
@@
STEAL-Net[astrid2021synthetic] 54.3 51.7
Jigsaw[wang2022video] 62.9 59.3
rGAN[lu2020few] 76.6 77.9*
MPN[lv2021learning] 78.9 73.8*
zxVAD[aich2023cross] 82.2 71.6*
Shibao et al.[gao2024scene] 86.2 78.7
ZS CLIP[radford2021learning] 62.3 60.9
ZS ImageBind[girdhar2023imagebind] 64.5 61.3
LLaVA-1.5[liu2024improved] 67.4 59.6
Video-ChatGPT[maaz2024video] 76.9 69.1